Consider a two-dimensional scatterplot representing the relationship between two continuous variables. If the correlation coefficient is -1, then:

a.

All points lie in a straight line with a slope of -1.

b.

All points lie in a straight line with an unknown negative slope.

c.

All points do not lie in a straight line but the best fitting regression line has a slope of -1.

d.

There is a strong positive relationship between the two variables.
Answer #1

Ans:

Option b is correct.

All points lie in a straight line with an unknown negative slope.

(as slope=r*sy/sx=-sy/sx)